# Room-temperature spin-lifetime anisotropy exceeding 60 in bilayer graphene spin valves proximity coupled to WSe2

### v1

@@ Abstract
A spin lifetime anisotropy between in-plane and out-of-plane spins in bilayer graphene (BLG) can be achieved by spin-orbit proximity coupling of graphene to transition metal dichalcogenides. This coupling reduces the in-plane spin lifetime due to proximity-induced spin scattering, while the out-of-plane spin lifetime remains largely unaffected. We show that at room temperature spin lifetime anisotropy exceeds 60 in a bilayer graphene lateral spin valve proximity coupled to WSe2. The out-of-plane spin lifetime of about 250 ps closely matches that of a BLG reference region not in contact with WSe2. In contrast, the estimated in-plane spin lifetime of less than 4 ps leads to a complete suppression of the in-plane spin signal at the ferromagnetic Co/MgO spin detector. The proximity coupling of WSe2 to BLG is particularly promising, as it does not compromise the charge carrier mobility within the graphene channel.
@@ I Introduction
A promising method to manipulate the spin transport in graphene-based spin-valve devices is to couple the graphene transport channel to transition metal dichalcogenides (TMDs) like WSe2, WS2, MoSe2, or MoS2 with their high spin-orbit coupling (SOC).[1, 2, 3, 4, 5] When proximity-coupled to graphene, the TMDs strong SOC imprints an effective out-of-plane spin orbit field on the graphene layer which enhances spin scattering for spins oriented in the graphene plane while leaving out-of-plane spins largely unaffected [6, 7]. This yields a pronounced spin lifetime anisotropy, defined as $\xi=\tau_{\perp}/\tau_{\parallel}$, between the out-of-plane $\tau_{\perp}$ and in-plane $\tau_{\parallel}$ spin relaxation times [8]. Theory has predicted that such proximity-induced spin–valley locking in graphene could produce giant anisotropy factors on the order of tens to hundreds [6].
In lateral spin-valve experiments on TMD/single-layer graphene (SLG) heterostructures utilizing MoS2, MoSe2, and WS2 spin-lifetime anisotropies of about $\xi\approx 10$ have been reported so far [9, 10, 11], whereas the theoretical expectation is at least one order of magnitude higher [6]. Furtheremore, the anisotropy in these studies was not sufficiently high enough to completely suppress the in-plane spin signal from reaching the detector electrode [9, 10]. Even though WSe2 has already been used in lateral spin valves, no anisotropy was determined or quantified that way [12, 13], but in weak antilocalization measurements on WSe2/SLG/hBN-heterostructures at temperatures below 2 K a lower bound of $\xi=20$ has been found [14].
@@
WSe2 is a substrate that leaves the charge carrier mobility largely unaffected in its graphene-based heterostructures [19, 20]. Therefore, WSe2 is expected to be particularly suitable for introducing high spin lifetime anisotropy into graphene-based spin valves through proximity coupling, while retaining its charge transport properties.
Here we report a room temperature spin-lifetime anisotropy with a lower bound of 60 obtained in a lateral spin-valve device consisting of bilayer graphene coupled to WSe2. The high anisotropy leads to a complete suppression of the in-plane spin signal at the ferromagnetic detector electrode in both in-plane Hanle and spin-valve measurements. Instead, the spin signal could be recovered in oblique Hanle measurements and x-Hanle measurements involving the out-of-plane spin component [21, 22, 23, 24]. Because of the full suppression of the in-plane component, the respective lifetime was estimated by numerical simulations of the angle-dependent, anisotropic spin precession data to be less than 4 ps.
II Results and Discussion
The device discussed here is built from exfoliated bilayer graphene on a Si++/SiO2 substrate, where one part of the BLG transport channel is proximity coupled to a multilayer WSe2 flake (see Figs. 1a and 1b). The WSe2 flake was placed across the BLG with a dry transfer technique using a PDMS-droplet [25] and is located in sample region C in Fig. 1b, further referred to as the ”WSe2 region”. Two uncovered regions labeled A and B in Fig. 1b serve as reference regions to characterize spin transport.
The spin-sensitive electrodes were fabricated from 3 nm magnesium oxide (MgO), serving as a tunnel barrier and 35 nm cobalt (Co) [26, 27], both deposited under ultra-high vacuum conditions [28]. The contacts have a width of 150-600 nm in the x-direction (see Fig. 1d for coordinate system). For such a contact geometry only a single magnetic domain is formed with its easy-axis along the y-direction [29]. A magnetic field in either the x-direction (in-plane perpendicular to the contacts) or in the z-direction (out-of-plane) slowly rotates the magnetization from the easy-axis with increasing field strength [30], which we accounted for in our data analysis by a Stoner-Wohlfarth model [31, 15, 10]. All measurements were performed at room temperature. For the non-local spin transport measurements a low-frequency lock-in technique was used with a frequency of $f=21\,\mathrm{Hz}$ and an AC current of $I_{\mathrm{AC}}$ = 10 $\mu$A added on a DC current of $I_{\mathrm{DC}}$ = 25 $\mu$[32].
@@ II.1 Spin transport in the reference regions
We first study the spin transport properties of the reference regions by performing spin-valve and Hanle measurements at room temperature (RT) at zero gate voltage near the charge neutrality point in a non-local configuration (Fig. 1a). For the spin-valve measurements, the magnetic field is changed in its magnitude along the y-direction (see Fig. 1c), i.e., at angle $\beta=0^{\circ}$ between the y and z-direction. Fig. 2a shows the measurement of region A. The spin-valve measurement (black curve) shows typical switching with a non-local resistance $\Delta R_{\text{nl}}$ of about 2.5 $\Omega$ (3.6 $\Omega$ for region B). From the spin precession measurements in Hanle configuration with a magnetic field perpendicular (i.e., $\beta=90^{\circ}$) to the sample plane (red curve in Fig. 2a), the spin lifetime $\tau_{s}$ and spin diffusion length $\lambda_{s}$ were determined with the steady-state solution of the Bloch-Torrey equation [34, 35]. To better distinguish the different measurement configurations, we will further refer to this as z-Hanle. The corresponding spin lifetime is 360 ps (280 ps for region B) with a spin diffusion length of $\lambda_{s}$=2.8 $\mu$m (1.7 $\mu$m for region B).
@@
The saturation resistance of the reference region $R_{\mathrm{nl}\beta}$ is normalized with $R_{\mathrm{nl}0}$ and plotted in Fig. 2d against $\cos^{2}(\beta)$. In an isotropic case ($\tau_{\perp}=\tau_{\parallel}$), a linear relationship versus $\cos^{2}\beta$ is expected as indicated by the straight line in Fig. 2d [9]. For anisotropies $\xi>1$ ($<1$), the curve would be above (below) the $\xi=1$ curve (black line). The fit to our data (red line) gives an anisotropy of $\xi=1.19$. According to Ref. [21] a value slightly above 1 can be explained by small out-of-plane spin-orbit fields that are caused by ripples or flexural distortions. However, in our sample we also observe that the oblique Hanle curves do not completely saturate even at higher magnetic fields as can be seen in the non-vanishing slope above 150 mT in all curves with $\beta\neq 0$ in Fig. 2b. This slope can be explained by a Hall signal, that is caused by a spatially inhomogeneous injection of charge carriers through pinholes in the tunnel barrier [36]. Since this Hall signal depends on the magnetic field direction (see [36]), it also varies with $\beta$ and therefore influences the previously described method for evaluating the spin lifetime anisotropy.
@@
To investigate this possible source of error with regard to the extracted spin lifetime anisotropy further, we performed x-Hanle measurements (Fig. 2c), in which the in-plane direction of the magnetic field does not result in any Hall voltages. In the isotropic case, the shape of the curve is expected to be identical to the z-Hanle curve, since the projection of the precessing spin ensemble on the y-direction (direction of the magnetization of the contacts) is identical regardless if the spins precess in the z-y- or x-y-plane. However, due to the shape anisotropy of the contacts, the magnetization of the contacts rotates significantly more when applying the magnetic field along the x-direction than along the z-direction [30]. The more the magnetization rotates towards the magnetic field direction, the larger the overall contribution of the injected spins (which are injected parallel to the magnetization) that is parallel to the external field. This non-precessing contribution of the injected spins results in an additonal background signal, that increases the more the magnitization is rotated away from the y-direction. This background is seen in a Fig. 2c as a steadily growing signal away from $B=0$ in both directions, while $R_{\mathrm{nl}}$ above $|B|>150\,\mathrm{mT}$ in the z-Hanle configuration approaches $R_{\mathrm{nl}}=0$ (see Fig. 2a).
We implemented a Stoner-Wohlfarth model [31, 15, 10, 37] in the steady-state solution of the Bloch-Torrey equation [35] to account for the rotation of the magnetization of the contacts in the x-Hanle measurement scheme. With the adapted fit (see dashed black line in Fig. 2c), we find a spin-lifetime of 350 ps in reference region A, which is within the fit error identical to the lifetime extracted from the z-Hanle measurement (360 ps). This implies an anisotropy factor of 1, as expected from graphene without any proximity coupling [15], in contrast to the value of 1.19 extracted from the oblique Hanle method. This indicates that charge-induced Hall signals must be considered if precise anistropy values are to be extracted from oblique Hanle measurements.
Figure 3: x-Hanle measurement and its numerical simulations in the WSe2 region. (a) x-Hanle measurement and best simulation result (green lines) with $\tau_{\perp}$=300 ps and $\xi=60$. (b) Variation of the spin-lifetime anisotropy $\xi$ for a constant $\tau_{\perp}$=300 ps. (c) Variation of $\tau_{\perp}$ for a constant $\xi=60$.
@@ II.2 WSe2 proximity-coupled region
Figure 4: Symmetrized and normalized oblique Hanle measurements of the WSe2-region and simulation curves. (a) Measurement set at different angles $\beta$ with corresponding simulations using the best match parameters from (b) and (c), i.e. $\tau_{\perp}$=250 ps and $\xi=60$. (b) Measurement at $\beta=40^{\circ}$ and simulated curves for which the spin-lifetime anisotropy $\xi=60$ was kept constant while varying the out-of-plane spin lifetime $\tau_{\perp}$. (c) Similarly to (b) $\tau_{\perp}$=250 ps is kept constant while varying $\xi$.
@@
II.3 Conclusions
In summary, we have demonstrated a large room temperature spin lifetime anisotropy ($\xi\geq 60$) in bilayer graphene spin valves proximity-coupled to WSe2. Through a combination of Hanle and oblique spin precession measurements and supported by numerical simulations, we determined that the out-of-plane spin lifetime in the BLG/WSe2 region is approximately 250–300 ps, whereas the in-plane spin lifetime is suppressed below 4 ps. The latter effectively eliminates transport of in-plane spins, allowing only the out-of-plane oriented spins to propagate to the spin detector. The achieved spin lifetime anisotropy of 60 is the highest at 300 K reported so far for graphene-based heterostructures.
We primarily attribute the large spin lifetime anisotropy observed in our study to the use of BLG, which is predicted to exhibit significantly higher anisotropy compared to SLG [3, 18]. We note, however, that the twist angle between the TMD and BLG might also have an impact on the anisotropy. Prior studies have demonstrated that the strength of the SO proximity effect depends on the twist angle [39, 40, 41, 42, 43]. In our sample, we estimate the twist angle between the flakes to be approximately 18, based on microscope image and the assumption that the flakes have ripped along the same crystallographic axes during exfoliation. This estimated angle falls in the regime of strong Valley-Zeeman coupling [39, 42] which favors large spin lifetime anisotropy. However, further studies are required to determine the dependence of twist angle on spin lifetime anisotropy.
@@ Acknowledgements.
Notes: The authors declare no competing financial interests.

### v2

@@ Abstract
A spin lifetime anisotropy between in-plane and out-of-plane spins in bilayer graphene (BLG) can be achieved by spin-orbit proximity coupling of graphene to transition metal dichalcogenides. This coupling reduces the in-plane spin lifetime due to proximity-induced spin scattering, while the out-of-plane spin lifetime remains largely unaffected. We show that at room temperature spin lifetime anisotropy exceeds 60 in a bilayer graphene lateral spin valve proximity coupled to WSe2. The out-of-plane spin lifetime of about 250 ps closely matches that of a BLG reference region not in contact with WSe2. In contrast, the estimated in-plane spin lifetime of less than 4 ps leads to a complete suppression of the in-plane spin signal measured at the ferromagnetic Co/MgO spin detector. The proximity coupling of WSe2 to BLG is particularly promising, as it does not compromise the charge carrier mobility within the graphene channel.
@@ I Introduction
A promising method to manipulate the spin transport in graphene-based spin-valve devices is to couple the graphene transport channel to a material with high spin-orbit coupling [1, 2, 3]. Such materials include transition metal dichalcogenides (TMDs) like WSe2, WS2, MoSe2, MoS2 or their alloys [4, 5, 6], but also materials like topological insulators or two-dimensional magnetic insulators, where the latter includes composites such as Cr2Ge2Te2 or MnPSe3 that can induce ferro- or antiferromagnetic proximity exchange coupling in graphene [7, 8, 9]. Due to their strong SOC, TMDs that are proximity-coupled to graphene imprint an effective out-of-plane spin orbit field on the graphene layer, which enhances spin scattering for spins oriented in the graphene plane while leaving out-of-plane spins largely unaffected [10, 11, 1, 4, 5, 7]. This yields a pronounced spin lifetime anisotropy, defined as $\xi=\tau_{\perp}/\tau_{\parallel}$, between the out-of-plane ($\tau_{\perp}$) and in-plane ($\tau_{\parallel}$) spin relaxation times [12].
In lateral spin-valve experiments on TMD/single-layer graphene (SLG) heterostructures utilizing MoS2, MoSe2, WS2, and PdSe2 spin-lifetime anisotropies of about $\xi\approx 10$ have been reported so far [13, 14, 15, 16], whereas the theoretical expectation is at least one order of magnitude higher [10]. Furtheremore, the anisotropy in these studies was not sufficiently high enough to completely suppress the in-plane spin signal from reaching the detector electrode [13, 14, 15, 16]. Even though WSe2 has already been used in lateral spin valves, no anisotropy was determined or quantified that way [17, 18], but in weak antilocalization measurements on WSe2/SLG/hBN-heterostructures at temperatures below 2 K a lower bound of $\xi=20$ has been found [19].
Furthermore, using bilayer graphene (BLG) instead of single layer graphene offers additional opportunities for graphene-based spintronic devices that are proximity coupled to TMDs. Electrical transport measurements in BLG devices that are in contact with TMDs on either one or both sides reveal that the proximity-induced SOC is layer-selective, primarily impacting the graphene layer in direct contact with the TMD [20, 21, 22]. Combined with the fact that the band structure of BLG can be tuned with out-of-plane electric fields to either open a band-gap [23] or to confine all charge carriers to only one of the two layers [24], enables device concepts such as spin-orbit valves or spin transistors [24, 4]. In this respect, a BLG spin-valve that was fully supported by a large WS2 flake has already shown a spin-lifetime anisotropy of about $\xi=40\dots 70$ albeit at low temperatures of 4 K [25]. However, much higher values up to $\xi=10\,000$ were theoretically predicted in BLG proximity-coupled to WSe2 [4]. At the same time, WSe2 is a substrate that leaves the charge carrier mobility largely unaffected in its graphene-based heterostructures [26, 27]. Therefore, WSe2 is expected to be particularly suitable for introducing high spin lifetime anisotropy into graphene-based spin valves through proximity coupling, while retaining its charge transport properties.
Here we report a room temperature spin-lifetime anisotropy with a lower bound of 60 obtained in a lateral spin-valve device consisting of bilayer graphene coupled to WSe2. The high anisotropy leads to a complete suppression of the in-plane spin signal measured at the ferromagnetic detector electrode in both in-plane Hanle and spin-valve measurements. Instead, the spin signal could be recovered in oblique Hanle measurements and x-Hanle measurements involving the out-of-plane spin component [28, 29, 30, 31]. Because of the full suppression of the in-plane component, the respective lifetime was estimated by numerical simulations of the angle-dependent, anisotropic spin precession data to be less than 4 ps.
II Results and Discussion
@@ II.1 Spin transport in the reference regions
[…]
@@
The saturation resistance of the reference region $R_{\mathrm{nl}\beta}$ is normalized with $R_{\mathrm{nl}0}$ and plotted in Fig. 2d against $\cos^{2}(\beta)$. In an isotropic case ($\tau_{\perp}=\tau_{\parallel}$), a linear relationship versus $\cos^{2}\beta$ is expected as indicated by the straight line in Fig. 2d [13]. For anisotropies $\xi>1$ (<1<1), the curve would be above (below) the $\xi=1$ curve (black line). The fit to our data (red line) gives an anisotropy of $\xi=1.19$. According to Ref. [28] a value slightly above 1 can be explained by small out-of-plane spin-orbit fields that are caused by ripples or flexural distortions. However, in our sample we also observe that the oblique Hanle curves do not completely saturate even at higher magnetic fields as can be seen in the non-vanishing slope above 150 mT in all curves with $\beta\neq 0$ in Fig. 2b. This slope can be explained by a Hall signal, that is caused by a spatially inhomogeneous injection of charge carriers through pinholes in the MgO barrier [44]. Since this Hall signal depends on the magnetic field direction (see [44]), it also varies with $\beta$ and therefore influences the previously described method for evaluating the spin lifetime anisotropy.
@@
To investigate this possible source of error with regard to the extracted spin lifetime anisotropy further, we performed x-Hanle measurements (Fig. 2c), in which the in-plane direction of the magnetic field does not result in any Hall voltages. In the isotropic case, the shape of the curve is expected to be identical to the z-Hanle curve, since the projection of the precessing spin ensemble on the y-direction (direction of the magnetization of the contacts) is identical regardless if the spins precess in the z-y- or x-y-plane. However, due to the shape anisotropy of the contacts, the magnetization of the contacts rotates significantly more when applying the magnetic field along the x-direction than along the z-direction [37]. The more the magnetization rotates towards the magnetic field direction, the larger the overall contribution of the injected spins (which are injected parallel to the magnetization) that is parallel to the external field. This non-precessing contribution of the injected spins results in an additional background signal, that increases the more the magnetization is rotated away from the y-direction. This background is seen in a Fig. 2c as a steadily growing signal away from $B=0$ in both directions, while $R_{\mathrm{nl}}$ above $|B|>150\,\mathrm{mT}$ in the z-Hanle configuration approaches $R_{\mathrm{nl}}=0$ (see Fig. 2a).
We implemented a Stoner-Wohlfarth model [38, 39, 14, 45] in the steady-state solution of the Bloch-Torrey equation [43] to account for the rotation of the magnetization of the contacts in the x-Hanle measurement scheme. With the adapted fit (see dashed black line in Fig. 2c), we find a spin-lifetime of 350 ps in reference region A, which is within the fit error identical to the lifetime extracted from the z-Hanle measurement (360 ps). This implies an anisotropy factor of 1, as expected from graphene without any proximity coupling [39], in contrast to the value of 1.19 extracted from the oblique Hanle method. This indicates that charge-induced Hall signals must be considered if precise anisotropy values are to be extracted from oblique Hanle measurements.
Figure 3: x-Hanle measurement and its numerical simulations in the WSe2 region. (a) x-Hanle measurement and best simulation result (green lines) with $\tau_{\perp}$=300 ps and $\xi=60$. (b) Variation of the spin-lifetime anisotropy $\xi$ for a constant $\tau_{\perp}$=300 ps. (c) Variation of $\tau_{\perp}$ for a constant $\xi=60$. The spin diffusion coefficient is set to 0.025 m2/s for the simulations.
@@ II.2 WSe2 proximity-coupled region
Figure 4: Symmetrized and normalized oblique Hanle measurements of the WSe2-region and simulation curves. (a) Measurement set at different angles $\beta$ with corresponding simulations using the best match parameters from (b) and (c), i.e. $\tau_{\perp}$=250 ps and $\xi=60$. (b) Measurement at $\beta=40^{\circ}$ and simulated curves for which the spin-lifetime anisotropy $\xi=60$ was kept constant while varying the out-of-plane spin lifetime $\tau_{\perp}$. (c) Similarly to (b) $\tau_{\perp}$=250 ps is kept constant while varying $\xi$. The spin diffusion coefficient is set to 0.025 m2/s for the simulations.
@@
II.3 Underestimation of the spin lifetime anisotropy
Although the extracted spin lifetime anisotropy of $\xi\geq 60$ is one of the highest values reported in literature so far, it remains well below the theoretical prediction of up to $\xi=10\,000$ in case of BLG proximity-coupled to WSe2 [4]. The most likely explanation for this is that we have neglected the impact of contacts and substrates on the measured spin lifetimes. It has been demonstrated that using the steady-state solution of the Bloch-Torrey equation [42, 43] without considering contact- or substrate-induced spin relaxation processes leads to an underestimation of the extracted spin lifetime [47, 48, 34, 49, 50, 46, 51, 52, 53, 54].
The measured total spin relaxation rate (i.e. the inverse of the measured spin lifetime) is the sum of contributions from spin relaxation rates due to the contacts, the SiO2 substrate and the proximity-induced out-of-plane spin orbit field caused by the TMD. Assuming that only the latter exhibits an anisotropy between out-of-plane and in-plane spin lifetimes, whereas the spin relaxation rates due to the SiO2 substrate and the MgO/Co contacts are the same for both spin orientations, we can make the following reasoning: In case of the in-plane spin lifetime, the complete suppression of the in-plane spin component implies that the proximity-induced out-of-plane spin orbit field completely dominates over other relaxation channels. The opposite is true for the out-of-plane spin lifetime, at it is very similar to the spin lifetime in the reference region, implying that substrate- and contact-induced relaxation rates dominate here.
At the same time, it was shown that spin lifetimes of up to 12 ns can be reached in graphene-based spin valve devices [55]. This implies that the determined out-of-plane spin lifetime of 250–300 ps could, in principle, be enhanced by at least a factor of 40, if substrate- and contact-induced relaxation processes could be diminished. Under such conditions, the corresponding spin lifetime anisotropy could potentially increase to $\xi=60\times 40=2\,400$, bringing it much closer to the theoretically expected value of $\xi=10\,000$.
Furthermore, we note that the twist angle between the TMD and BLG might also have an impact on the anisotropy. Prior studies have demonstrated that the strength of the SO proximity effect depends on the twist angle [56, 57, 58, 59, 60, 9, 61, 62]. In our sample, we estimate the twist angle between the flakes to be approximately 18, based on microscope images and the assumption that the flakes have ripped along the same crystallographic axes during exfoliation. This estimated angle falls in the regime of strong Valley-Zeeman coupling [56, 59] which favors large spin lifetime anisotropies. However, further studies are required to determine the dependence of twist angle on spin lifetime anisotropy.
II.4 Conclusions
In summary, we have demonstrated a large room temperature spin lifetime anisotropy of $\xi\geq 60$ in bilayer graphene spin valves proximity-coupled to WSe2. Through a combination of Hanle and oblique spin precession measurements and supported by numerical simulations, we determined that the out-of-plane spin lifetime in the BLG/WSe2 region is approximately 250–300 ps, whereas the in-plane spin lifetime is suppressed below 4 ps. The latter effectively eliminates transport of in-plane spins, allowing only the out-of-plane oriented spins to propagate in the direction of the spin detector. The achieved spin lifetime anisotropy of 60 is the highest at 300 K reported so far for graphene-based heterostructures. At the same time, the determined anisotropy factor represents a conservative lower bound of what can be achieved in a BLG/WSe2 heterostructure, as substrate- and contact-induced spin relaxation processes have likely limited the measured out-of-plane spin lifetime. We attribute the large spin lifetime anisotropy to the use of BLG, which is predicted to exhibit significantly higher anisotropy compared to SLG [4, 25].
@@ Acknowledgements.
Notes: The data that support the findings of this study are openly available at https://doi.org/10.5281/zenodo.15609970.